Consider an asset that costs $725,000 and is depreciated straight-line to zero over its nine-year tax life. The asset is to be used in a five-year project; at the end of the project, the asset can be sold for $161,000. If the relevant tax rate is 23 percent, what is the aftertax cash flow from the sale of this asset?
Answer #1

Annual depreciation=(Cost-Accumulated Depreciation)/Useful Life

(725000/9)=$80555.56/year(Approx)

Hence book value as on date of sale=Cost-Accumulated depreciation

=725000-(80555.56*5)=$322,222.22(Approx)

Hence loss on sale=(322,222.22-161000)=$161,222.22(Approx)

Hence cash flow=Sale proceeds+(Tax rate*loss on sale)

=161000+(161,222.22*0.23)

$198081.11(Approx).